# Galois Action and Localization in Number Fields

### v1

@@ Abstract.
For a Galois number field $K$, the Galois group $\text{Gal}(K/\mathbb{Q})$ acts on the class group $Cl_{K}$ in a very natural way: $\sigma\cdot[I]=[\sigma(I)]$ for any $\sigma\in\text{Gal}(K/\mathbb{Q})$, $I]\in Cl_{K}$. In this paper, we will explore how the unique properties of this group action work together to elucidate the relationship between these two groups. While previous work on this problem has focused on representation theory, we take a direct approach to some classical and new problems. The paper concludes with an exploration of the class groups of localizations of the ring of integers $\mathcal{O}_{K}$. These turn out to be powerful tools for understanding $Cl_{K}$ and overrings of $\mathcal{O}_{K}$.
@@ 1. Introduction and Notation
Throughout this paper $K$ will denote a Galois number field with Galois group $G:=\text{Gal}(K/\mathbb{Q})$, and $\mathcal{O}_{K}$ its ring of integers with class group $Cl_{K}$ and class number $h_{K}=|Cl_{K}|$. The Galois group of $K$ acts on the class group in a very natural way. For any $\sigma\in G$ and $I]\in Cl_{K}$, we can define the action $\sigma\cdot[I]=[\sigma(I)]$. This is a well-defined group action which provides us with some tools for characterization the relationship between the two groups. More than this, we also have the peculiar property that $\sigma\cdot([I][J])=(\sigma\cdot[I])(\sigma\cdot[J])$ for any $\sigma\in G$ and $I],[J]\in Cl_{K}$. This induces a map from $G$ to $\text{Aut}(Cl_{K})$ given by $\sigma\mapsto\bar{\sigma}$ where $\bar{\sigma}([I])=[\sigma(I)]$. Finally, we have the “norm property” that $\prod_{\sigma\in G}\sigma\cdot[I]=\left[\prod_{\sigma\in G}\sigma(I)\right]=\text{Prin}(\mathcal{O}_{K})$ for any $I]\in G$—for details, see [12]. Note, in this paper we will study the case where $K$ is Galois over $\mathbb{Q}$ with the knowledge that the theorems herein apply also to the case when $K/L$ is a Galois extension of number fields with $\mathcal{O}_{L}$ a principal ideal domain (PID). Now, these characteristics of the action motivate the following definition.
@@
In the following section, we will take a closer look at the conditions of Definition 1.1 and their implications for norm-like actions in general. In Section 3, we explore how conditions 1-2, together with 4, place restrictions of the structure of $Cl_{K}$ for a Galois number field of given degree. In Section 4, we focus on applications of condition 3 to the inverse class group problem. In particular, which number fields $K$ can have a given abelian group as their class group. Finally, in Section 5, the authors determine the structure of the class group for localizations of the form $\mathcal{O}_{K}[\frac{1}{x}]$ which ultimately strengthens many of the previously developed results and leads to some interesting observations about rings of integers.
@@ 2. Norm-Like Group Actions
Conditions $1$ and $2$ are those of a typical group action. The definition diverges with condition $3$ which ensures the action of each element $g$ on $A$ behaves like a homomorphism. In fact, each element of $G$ can be identified with an endomorphism of $A$ via the map $g\mapsto\phi$ where $\phi(a)=g\cdot a$. If $A$ is finite, in the case where $g\cdot a=e_{A}\implies a=e_{A}$, this map defines a group homomorphism from $G$ to $\text{Aut}(A)$, the kernel of which is the set $\{g\in G\,|\,g\cdot a=a\,\forall\,a\in A\}$. The implications of this are closely considered in Section 4. We also note that condition $3$ guarantees $g\cdot e_{A}=e_{A}$ for all $g\in G$. Thus, a norm-like group action will be transitive if and only if $A$ is trivial.
@@
Now, let $A$ be an arbitrary abelian group, and $G=\{id,-id\}\leqslant\text{Aut}(A)$ where $id$ is the identity, and id(a)=a-id(a)=-a for all $a\in A$, which is a automorphism just in case $A$ is abelian. It is easy to verify that the natural action $g\cdot a=g(a)$ is norm-like. This case has been of particular interest over the last three years as the study of weighted zero-sum sequences has become popular in connection with normset factorization (see [1] and [7]). In particular, $\{id,-id\}$ is the subgroup induced by the action of $\text{Gal}(K/\mathbb{Q})$ on $Cl_{K}$ when $K$ is quadratic. As we have seen, $|G|=2$ places no restriction on $A$. As we will see throughout this paper, this is a very unique case. Even $|G|=3$ is much more restrictive and the relevant arithmetic far more complicated.
@@ 3. Direct Consequences of the Group Action
Let $K$ be a Galois number field of degree $p^{r}$. Then, $h_{K}\equiv 0\text{ or }1\text{ mod }p$.
@@
We now turn our attention to some results on Galois number fields of odd degree.
@@
Let $K$ be a Galois number field of odd degree. Then, $Cl_{K}$ cannot have a unique involution.
@@
Assume for the purpose of contradiction that $Cl_{K}$ has a unique element of order 2, call it $[J]$. Thus, as each $\sigma\in G$ is an automorphism, we must have $\sigma\cdot[J]=[\sigma(J)]=[J]$ for all $\sigma\in G$. Hence, $G$ acts trivially on $[J]$, so by Lemma 3.1 the order of $[J]$ divides $|G|$. However, this is a contradiction as $[J]$ has order $2$ and $|G|$ is odd. ∎
@@
Note that this result also follows directly from Theorem 3.3. We conclude this section with a specific application of these results.
@@
Recall that the action of $G=\text{Gal}(K/\mathbb{Q})$ on $Cl_{K}$ induces a homomorphism from $G$ to $\text{Aut}(Cl_{K})$ given by $\sigma\overset{\psi}{\mapsto}\phi$ where $\phi([I])=\sigma\cdot[I]=[\sigma(I)]$. In this section, we will see how this map can be used to further restrict the structure of $Cl_{K}$ for a Galois number field $K$. Conversely, we will use this tool to help answer questions related to the inverse class group problem. This terminology is used in the spirit of the inverse Galois problem. In particular, the inverse class group problem asks if an arbitrary abelian group can be realized as the class group of a ring of integers. This question was famously answered in the positive for the more general class of Dedekind domains by Claborn in [4]. In the case of imaginary quadratics, the answer is known to be no. Certain groups can be shown by brute force not to appear, the smallest among them being $(\mathbb{Z}/3\mathbb{Z})^{3}$. For details, the reader is encouraged to see (CITE). The question remains open for rings of integers in general. In this section, for a given finite abelian group $A$, we will explore which rings of integers may admit $Cl_{K}\cong A$. We begin with a direct proof of another result from [5] which we will return to with some new found tools in section 4.
@@
Say $\phi(1)=a$, then we must have $\phi(x)=a\cdot x$ for any $x\in\mathbb{Z}/p\mathbb{Z}$. This also implies that $a$ is a primitive root mod $p$. Now, any $x\in\mathbb{Z}/p\mathbb{Z}$ is contained in a (p1)/n(p-1)/n cycle in $\phi^{n}$. Thus, the cycle is of the form $(x\,\,\,xa^{n}\,\,\,xa^{2n}\dotsm xa^{(\frac{p-1}{n}-1)n})$. Taking the sum, we get
@@ 5. Localizations of Rings of Integers
In this section, we will use localization to strengthen some of the results developed previously. First, recall that any overring—in particular, any localization—of a Dedekind domain remains Dedekind (for further information see CITE). Intuitively, as a localization of $R$ is attained by turning a set of elements $S\subseteq R$ into units, we expect the class group of $R_{S}$ to be smaller than that of $R$ as those non-principal ideals $I\subseteq R$ for which $I\cap S\neq\emptyset$ will become principal. As we will demonstrate, “small” localizations will tend to lead to small changes in the class group. In the context of rings of algebraic integers, this will be useful in further restricting the possible structure of $Cl_{K}$.
@@
Note, $D[\frac{1}{x}]$ is the localization of $D$ at the multiplicative set $\{x^{n},\,n\in\mathbb{N}\}$.
@@
The remainder of this section will be dedicated to leveraging this fact to extend the techniques and theorems developed in sections 2 and 3. Our hope is that the Galois group of $K$ will act on the class group of $\mathcal{O}_{K}[\frac{1}{x}]$ in the same manner as $Cl_{K}$ which would allow us to place the same restrictions on the homomorphic images of the class group—further constraining the structure of $Cl_{K}$. Unfortunately, the action $\sigma\cdot[\mathfrak{P}]=[\sigma(\mathfrak{P})]$ for $\sigma\in G$ and $\mathfrak{P}]\in Cl(\mathcal{O}_{K}[\frac{1}{x}])$ will not be well-defined in general as $\sigma(\frac{1}{x})=\sigma(x)^{-1}$ need not be in $\mathcal{O}_{K}[\frac{1}{x}]$ in general. However, in the case that $x\in\mathbb{Z}$, we avoid this issue as we will have $\sigma(\frac{1}{x})=\frac{1}{x}$ for all $\sigma\in G$.
@@
For a Galois number field $K$ and $x\in\mathcal{O}_{K}$ a nonzero nonunit, the group action $\sigma\cdot[\mathfrak{P}]=[\sigma(\mathfrak{P})]$ for $\sigma\in\text{Gal}(F/\mathbb{Q}),[\mathfrak{P}]\in Cl\left(\mathcal{O}_{K}[\frac{1}{N(x)}]\right)$ is a Galois action.
@@
It is sufficiently clear that conditions 1-3 from Definition 1.1 hold. Now, let $\mathfrak{P}]\in Cl(\mathcal{O}_{K}[\frac{1}{N(x)}])$ where $\mathfrak{P}$ is a prime ideal representative. If we write $\mathfrak{p}=\mathfrak{P}\mathbin{\cap}\mathcal{O}_{K}$, we have $\left(\prod_{\sigma\in G}\sigma(\mathfrak{p})\right)\mathcal{O}_{K}[\frac{1}{N(x)}]=\prod_{\sigma\in G}\sigma(\mathfrak{P})$, so the norm property (4) is also inherited from our original group action. ∎
@@
This was proven for abelian number fields by Fröhlich ([8]) and is a notably more restrictive condition. For example, observe that while Theorem 3.2 would not preclude an extension of degree $K:\mathbb{Q}]=3^{r}$ from having $h_{k}=55$, Corollary 5.7 shows this is impossible. In fact, such an extension cannot have $h_{K}=5n$ for any $n\in\mathbb{N}$ not divisible by $5$.
We conclude with a Jordan-Hölder like result which follows from Theorem 5.1.
@@
Let $D$ be a Dedekind domain with finite class group $Cl(D)$ such that every class contains a prime ideal and $Cl(D)$. Let $R$ be an overring of $D$ which is minimal with respect to being a PID. Then there exists a finite sequence of adjacent domains $D:=D_{0}\subseteq D_{1}\subseteq D_{2}\subseteq\dotsm\subseteq D_{r}:=R$.
@@
Because $Cl(D)$ is abelian, we may choose an arbitrary element $\mathfrak{q}]\in Cl(D)$ of prime order $p$ where $\mathfrak{q}$ is a prime ideal representative of the class. Allowing $(a)=\mathfrak{q}^{p}$, by Theorem 5.1, $D_{1}:=D[\frac{1}{a}]$ has class group isomorphic to $Cl(D)/\left<[\mathfrak{q}]\right$>. We now show that $D\subseteq D[\frac{1}{a}]$ is a minimal ring extension.
Assume there exists some intermediate ring $D\subsetneq A\subseteq D[\frac{1}{a}]$. It is well known (see [9]) that
$A=\underset{\mathfrak{p}A\neq A}{\bigcap}D_{\mathfrak{p}}$
where the $\mathfrak{p}$ range over $\text{Spec}(D)$. Similarly,
$D=\underset{\mathfrak{m}\in\text{MaxSpec}(D)}{\bigcap}D_{\mathfrak{m}}=\underset{\mathfrak{p}\in\text{Spec}(D)}{\bigcap}D_{\mathfrak{p}}$
with the second equality due to $D$ being $1$-dimensional. As $A\neq D$, the above tells us there must exist some prime ideal $\mathfrak{p}^{\prime}\subseteq D$ such that $\mathfrak{p}^{\prime}A=A$ which implies some $d\in D$ becomes a unit in $A\subseteq D[\frac{1}{a}]$. Now, as $\mathfrak{p}]^{q}$ forms a minimal 0-sequence, we must have $a$ irreducible in $D$. Thus the saturation of $\{a^{n}\,|\,n\in\mathbb{N}\}$ in $D$ is simply $\{a^{n}\,|\,n\in\mathbb{N}\}$. Hence, $\frac{1}{a^{k}}\in A$ for some $k\in\mathbb{N}$ which implies $\frac{1}{a}\in A$, so we must have $A=D[\frac{1}{a}]$ as desired. Finally, as $Cl(D)$ is finite, we know that after a finite number of steps, we must have $Cl(D_{r})=1$, and thus $D_{r}$ is a PID.
Porism 5.9.
@@
Let $D$ be a Dedekind domain with torsion class group and $a\in D$ irreducible. Then $D\subseteq D[\frac{1}{a}]$ is a minimal ring extension.

### v2

@@ Abstract.
For a Galois number field $K$, the Galois group $\text{Gal}(K/\mathbb{Q})$ acts on the class group $Cl_{K}$ in a very natural way: $\sigma\cdot[I]=[\sigma(I)]$ for any $\sigma\in\text{Gal}(K/\mathbb{Q})$, $I]\in Cl_{K}$. In this paper, we will explore how the unique properties of this group action work together to elucidate the relationship between these two groups. While previous work on this problem has focused on representation theory, we take a direct approach to some classical and new problems. The paper concludes with an exploration of the class groups of overrings of the ring of integers $\mathcal{O}_{K}$. These turn out to be powerful tools for understanding $Cl_{K}$ and overrings of $\mathcal{O}_{K}$.
@@ 1. Introduction and Notation
Throughout this paper $K$ will denote a Galois number field with Galois group $G:=\text{Gal}(K/\mathbb{Q})$, and $\mathcal{O}_{K}$ its ring of integers with class group $Cl_{K}$ and class number $h_{K}=|Cl_{K}|$. The Galois group of $K$ acts on the class group in a very natural way. For any $\sigma\in G$ and $I]\in Cl_{K}$, we can define the action $\sigma\cdot[I]=[\sigma(I)]$. This is a well-defined group action which provides us with some tools for characterization the relationship between the two groups. More than this, we also have the peculiar property that $\sigma\cdot([I][J])=(\sigma\cdot[I])(\sigma\cdot[J])$ for any $\sigma\in G$ and $I],[J]\in Cl_{K}$. This induces a map from $G$ to $\text{Aut}(Cl_{K})$ given by $\sigma\mapsto\bar{\sigma}$ where $\bar{\sigma}([I])=[\sigma(I)]$. Finally, we have the “norm property” that $\prod_{\sigma\in G}\sigma\cdot[I]=\left[\prod_{\sigma\in G}\sigma(I)\right]=\text{Prin}(\mathcal{O}_{K})$ for any $I]\in Cl_{K}$—for details, see [13]. Note, in this paper we will study the case where $K$ is Galois over $\mathbb{Q}$ with the knowledge that the theorems herein apply also to the case when $K/L$ is a Galois extension of number fields with $\mathcal{O}_{L}$ a principal ideal domain (PID). Now, these characteristics of the action motivate the following definition.
@@
In the following section, we will take a closer look at the conditions of Definition 1.1 and their implications for norm-like actions in general. In Section 3, we explore how conditions 1-2 together with 4 place restrictions of the structure of $Cl_{K}$ for a Galois number field of given degree. In Section 4, we focus on applications of condition 3 to the inverse class group problem. In particular, which number fields $K$ can have a given abelian group as their class group. In Section 5, the authors determine the structure of the class group for localizations of the form $\mathcal{O}_{K}[\frac{1}{x}]$ which ultimately strengthens many of the previously developed results and leads to some interesting observations about rings of integers. Finally, Section 6 concludes with a more general exploration of overrings of Dedekind domains, culminating with a characterization of those overrings whose class groups admit a norm-like group action.
@@ 2. Norm-Like Group Actions
Conditions $1$ and $2$ are those of a typical group action. The definition diverges with condition $3$ which ensures the action of each element $g$ on $A$ behaves like a homomorphism. In fact, each element of $G$ can be identified with an endomorphism of $A$ via the map $g\mapsto\phi$ where $\phi(a)=g\cdot a$. If $A$ is finite, in the case where $g\cdot a=e_{A}\implies a=e_{A}$, this map defines a group homomorphism from $G$ to $\text{Aut}(A)$, the kernel of which is the set $\{g\in G\,|\,g\cdot a=a\,\forall\,a\in A\}$. Note, this is precisely the kernel of the group action. The implications of this are closely considered in Section 4. We also note that condition $3$ guarantees $g\cdot e_{A}=e_{A}$ for all $g\in G$. Thus, a norm-like group action will be transitive if and only if $A$ is trivial.
@@
Now, let $A$ be an arbitrary abelian group, and $G=\{id,-id\}\leqslant\text{Aut}(A)$ where $id$ is the identity, and id(a)=a-id(a)=-a for all $a\in A$ which is a automorphism just in case $A$ is abelian. It is easy to verify that the natural action $g\cdot a=g(a)$ is norm-like. This case has been of particular interest over the last three years as the study of weighted zero sum sequences has become popular in connection with normset factorization (see [1] and [8]). In particular, $\{id,-id\}$ is the subgroup induced by the action of $\text{Gal}(K/\mathbb{Q})$ on $Cl_{K}$ when $K$ is quadratic. As we have seen, $|G|=2$ places no restriction on $A$. As we will see throughout this paper, this is a very unique case. Even the $|G|=3$ case is much more restrictive and the relevant arithmetic far more complicated.
Example 2.1.
Cubic Example: $G\cong\mathbb{Z}/3\mathbb{Z}$ acting on $A\cong\mathbb{Z}/7\mathbb{Z}$ Also, any of the subgroups of $SL_{2}(\mathbb{Z}/2\mathbb{Z})$ (order $6$) acting on $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ seem to work. There might be something here. Probably want a non-example in the cubic case as well.Is this unique in the sense that the action of all subgroups of $\text{Aut}(A)$ work via the natural group action.
@@ 3. Direct Consequences of the Group Action
Let $K$ be a Galois number field of degree $p^{r}$ where $p$ is prime. Then, $h_{K}\equiv 0\text{ or }1\text{ mod }p$.
@@
The question of whether every finite abelian group can be realized as the class group of a quadratic ring of integers is a famous open problem. We will consider some related questions in the next section. For now, we now turn our attention to some results on Galois number fields of odd degree.
@@
Let $K$ be a Galois number field of odd degree. Then, $Cl_{K}\ncong\mathbb{Z}/2n\mathbb{Z}$ for any $n\in\mathbb{N}$.
@@
Assume for the purpose of contradiction that $Cl_{K}\cong\mathbb{Z}/2n\mathbb{Z}$. Then, $Cl_{K}$ has a unique element of order 2, call it $[J]$. Thus, as each $\sigma\in G$ is an automorphism, we must have $\sigma\cdot[J]=[\sigma(J)]=[J]$ for all $\sigma\in G$. Hence, $G$ acts trivially on $[J]$, so by Lemma 3.1 the order of $[J]$ divides $|G|$. However, this is a contradiction as $[J]$ has order $2$ and $|G|$ is odd. ∎
@@
Note that this result also follows directly from Theorem 3.3. We conclude this section with a specific application of these results to factorization in cyclotomic number fields. These number fields in particular played a central role in the development of modern factorization theory. In his attempt to prove Fermat’s Last Theorem, Ernst Kummer erroneously assumed that the cyclotomic ring of integers $\mathbb{Z}[\xi_{p}]$ was a unique factorization domain for any primitive $p^{th}$ root of unity. It was Dedekind’s remarkable observation that while unique factorization for elements was lost, unique factorization of ideals was retained for all rings of integers. Thus, we are able to turn questions about the factorization of elements in rings of integers into questions about ideals and, in turn, the class group. Being Galois in general, the cyclotomic number fields provide an apt setting for the applications of the results we have developed in this section.
@@
Recall that the action of $G=\text{Gal}(K/\mathbb{Q})$ on $Cl_{K}$ induces a homomorphism from $G$ to $\text{Aut}(Cl_{K})$ given by $\sigma\overset{\psi}{\mapsto}\phi$ where $\phi([I])=\sigma\cdot[I]=[\sigma(I)]$. In this section, we will see how this map can be used to further restrict the structure of $Cl_{K}$ for a Galois number field $K$. Conversely, we will use this tool to help answer questions related to the inverse class group problem. This terminology is used in the spirit of the inverse Galois problem. In particular, the inverse class group problem asks if an arbitrary abelian group can be realized as the class group of a ring of integers. This question was famously answered in the positive for the more general class of Dedekind domains by Claborn in [5]. In the case of imaginary quadratics, the answer is known to be no. Certain groups can be shown by brute force not to appear, the smallest among them being $(\mathbb{Z}/3\mathbb{Z})^{3}$. The question remains open for rings of integers in general. In this section, for a given finite abelian group $A$, we will explore which rings of integers may admit $Cl_{K}\cong A$. We begin with a direct proof of another result from [6] which we will return to with some new found tools in section 4.
@@
Say $\phi(1)=a$, then we must have $\phi(x)=a\cdot x$ for any $x\in\mathbb{Z}/p\mathbb{Z}$. As $\phi$ generates $\text{Aut}(\mathbb{Z}/p\mathbb{Z})$, this also implies that $a$ is a primitive root mod $p$. Now, any $x\in\mathbb{Z}/p\mathbb{Z}$ is contained in a (p1)/n(p-1)/n cycle in $\phi^{n}$. Thus, the cycle is of the form $(x\,\,\,xa^{n}\,\,\,xa^{2n}\dotsm xa^{(\frac{p-1}{n}-1)n})$. Taking the sum, we get
@@ 5. Localizations of Rings of Integers
In this section, we will use localization to strengthen some of the results developed previously. First, recall that any overring—in particular, any localization—of a Dedekind domain remains Dedekind. Intuitively, as a localization of $R$ is attained by turning a set of elements $S\subseteq R$ into units, we expect the class group of $R_{S}$ to be smaller than that of $R$ as those non-principal ideals $I\subseteq R$ for which $I\cap S\neq\emptyset$ will become principal. As we will demonstrate, “small” localizations will tend to lead to small changes in the class group. In the context of rings of algebraic integers, this will be useful in further restricting the possible structure of $Cl_{K}$.
@@
Note, $D[\frac{1}{x}]$ is the localization of $D$ at the multiplicative set $\{x^{n},\,n\in\mathbb{N}\}$ which is the complement of the set-theoretic union of all prime ideals not containing $x$.
@@
The remainder of this section will be dedicated to leveraging this fact to extend the techniques and theorems developed in sections 2 and 3. Our hope is that the Galois group of $K$ will act on the class group of $\mathcal{O}_{K}[\frac{1}{x}]$ in the same manner as $\text{Cl}_{K}$ which would allow us to place the same restrictions on the homomorphic images of the class group—further constraining the structure of $Cl_{K}$. Unfortunately, the action $\sigma\cdot[\mathfrak{P}]=[\sigma(\mathfrak{P})]$ for $\sigma\in G$ and $\mathfrak{P}]\in Cl(\mathcal{O}_{K}[\frac{1}{x}])$ will not be well-defined in general as $\sigma(\frac{1}{x})=\sigma(x)^{-1}$ need not be in $\mathcal{O}_{K}[\frac{1}{x}]$ in general. However, in the case that $x\in\mathbb{Z}$, we avoid this issue as we will have $\sigma(\frac{1}{x})=\frac{1}{x}$ for all $\sigma\in G$.
@@
For a Galois number field $K$ and $x\in\mathcal{O}_{K}$ a nonzero nonunit, the group action $\sigma\cdot[\mathfrak{P}]=[\sigma(\mathfrak{P})]$ for $\sigma\in\text{Gal}(F/\mathbb{Q}),[\mathfrak{P}]\in Cl\left(\mathcal{O}_{K}[\frac{1}{N(x)}]\right)$ is a norm-like action.
@@
It is sufficiently clear that conditions 1-3 from Definition 1.1 hold. Now, let $\mathfrak{P}]\in Cl(\mathcal{O}_{K}[\frac{1}{N(x)}])$ where $\mathfrak{P}$ is a prime ideal representative. If we write $\mathfrak{p}=\mathfrak{P}\cap\mathcal{O}_{K}$, we have $\left(\prod_{\sigma\in G}\sigma(\mathfrak{p})\right)\mathcal{O}_{K}[\frac{1}{N(x)}]=\prod_{\sigma\in G}\sigma(\mathfrak{P})$, so the norm property (4) is also inherited from our original group action. ∎
@@
This was proven for abelian number fields by Fröhlich ([9]) which is a notably more restrictive condition. For example, observe that while Theorem 3.2 would not preclude an extension of degree $K:\mathbb{Q}]=3^{r}$ from having $h_{k}=55$, Corollary 5.7 shows this is impossible. In fact, such an extension cannot have $h_{K}=5n$ for any $n\in\mathbb{N}$ not divisible by $5$.
6. Overrings of Dedekind Domains
We conclude with a general consideration of overrings of Dedekind domains with eye toward developing a general criterion for when the class group of an overring of $\mathcal{O}_{K}$ is a Galois module via the action we have been studying. In particular, for an overring $\mathcal{O}_{K}\subseteq R\subseteq K$, we would like to know when the Galois group preserves fractional ideals of $R$ and thus admits the same norm-like group action on $Cl(R)$. We saw in Section 5 that even in the case of simple extensions, this may fail. Consider the following example.
Example 6.1.
Let $\mathcal{O}_{K}=\mathbb{Z}[\sqrt{-5}]$ and consider the overring $\mathcal{O}_{K}\left[\frac{1}{1+\sqrt{-5}}\right]$. Now, $\text{Gal}(K/\mathbb{Q})=\{\mathbf{1},\sigma\}$ where $\sigma$ represents complex conjugation. Note that $\sigma(R)$ fails to be an $R$-module as $\frac{1}{1+\sqrt{-5}}\cdot\sigma\left(\frac{1}{1+\sqrt{-5}}\right)=\frac{1}{1+\sqrt{-5}}\cdot\frac{1}{1-\sqrt{-5}}=\frac{1}{6}\notin\sigma(R)$.
The following is a classical result which characterizes all overrings of Dedekind domains with torsion class group. For details, the reader is encouraged to see [10].
@@ 6. Overrings of Dedekind Domains
Let $D$ be a Dedekind domain with torsion class group and $R$ an overring of $D$. Then
$R=\underset{\mathfrak{p}R\neq R}{\bigcap}D_{\mathfrak{p}}$
where the $\mathfrak{p}$ range over $\text{Spec}(D)$.
From this, we can quickly derive the following theorem of which we will make great use.
Theorem 6.3.
Let $D$ be a Dedekind domain with torsion class group. There exists a 111-1 correspondence between overrings of $D$ and nonempty subsets of $\text{Spec}(D)$ given by
$S\longleftrightarrow\underset{\mathfrak{p}\in S}{\bigcap}D_{\mathfrak{p}}.$
@@
By Theorem 6.2, every overring of $D$ can be expressed in the form $\underset{\mathfrak{p}\in S}{\bigcap}D_{\mathfrak{p}}$ for some $S\subseteq\text{Spec}(D)$. Thus, it suffices to show that $\underset{\mathfrak{p}\in S_{1}}{\bigcap}D_{\mathfrak{p}}\neq\underset{\mathfrak{p}\in S_{2}}{\bigcap}D_{\mathfrak{p}}$ for subsets $S_{1},S_{2}\subseteq\text{Spec}(D)$ whenever $S_{1}\neq S_{2}$. Without loss of generality, assume $\mathfrak{q}\in S_{1}\backslash S_{2}.$ Recall that $Cl(D)$ is torsion, so $|[\mathfrak{q}]|:=n$ is finite. Thus, $\mathfrak{q}^{n}=(q)$ for some $q\in D$. Now, $\mathfrak{q}$ is the unique prime ideal containing $q$, so $\frac{1}{q}\in\underset{\mathfrak{p}\in S_{2}}{\bigcap}D_{\mathfrak{p}}\backslash\underset{\mathfrak{p}\in S_{1}}{\bigcap}D_{\mathfrak{p}}$. ∎
Corollary 6.4.
@@
Let $D$ be a Dedekind domain with torsion class group. Then $D\subseteq D[\frac{1}{a}]$ is a minimal ring extension if and only if $aD=\mathfrak{p}^{k}$ for some $k\in\mathbb{N}$.
Proof.
We prove the forward implication by contraposition. Assume there are distinct primes $\mathfrak{p},\mathfrak{q}\in\text{Spec}(D)$ which divide $aD$. Then, $a\in\mathfrak{p},\mathfrak{q}$, and hence $\mathfrak{p}D[\frac{1}{a}]=\mathfrak{q}D[\frac{1}{a}]=D$. Thus, by Corollary 6.3, there is an overring properly between $D$ and $D[\frac{1}{a}]$.
Now, for the reverse implication, assume $aD=\mathfrak{p}^{k}$ and $\mathfrak{q}D[\frac{1}{a}]=D[\frac{1}{a}]$ for some prime $\mathfrak{q}\in\operatorname{Spec}(D)$. Because $D=\mathfrak{p}^{k}, aa$ is strongly irreducible in $D$, and thus the saturation of $\{a^{n}\,|\,n\in\mathbb{N}\}$ in $D$ is simply $\{a^{n}\,|\,n\in\mathbb{N}\}$. Hence, $a^{k}\in\mathfrak{q}$ for some $k\in\mathbb{N}$ which implies $a\in\mathfrak{q}$. However, $\mathfrak{p}$ is the unique prime ideal containing $a$, so $\mathfrak{p}=\mathfrak{q}$. Therefore, $\mathfrak{p}$ is the unique prime ideal of $D$ such that $\mathfrak{p}D[\frac{1}{a}]=D[\frac{1}{a}]$. By the correspondence in Corollary 6.3, we have the $D\subseteq D[\frac{1}{a}]$ must be a minimal extension.
By Theorem 6.3, we know that any adjacent domain $D\subseteq D^{\prime}$, there is a unique prime $\mathfrak{p}\in\text{Spec}(D)$ such that $\mathfrak{p}D^{\prime}=D^{\prime}$. What is more, any adjacent domain is a simple extension, and this $\mathfrak{p}$ determines the $\alpha$ such that $D^{\prime}=D[\alpha]$ as $\alpha=\frac{1}{a}$ where $aD=\mathfrak{p}^{n}$ where $|\mathfrak{p}|=n$. From these observations, we have the following Jordan-Hölder like result on overrings of Dedekind domains with finite class group.
Theorem 6.5.
Let $D$ be a Dedekind domain with finite group $Cl(D)$ such that every class contains a prime ideal. Let $R$ be an overring of $D$ which is minimal with respect to being a PID. If $D:=D_{0}\subseteq D_{1}\subseteq D_{2}\subseteq\dotsm\subseteq D_{r_{1}}:=R$ and $D:=D_{0}^{\prime}\subseteq D_{1}^{\prime}\subseteq D_{2}^{\prime}\subseteq\dotsm\subseteq D_{r_{2}}^{\prime}:=R$ are sequences of adjacent domains, then $r_{1}=r_{2}$ and the sequence of $\alpha_{i}\in R$ such that $D_{i+1}=D_{i}[\alpha_{i}]$ is unique up to reordering. What is more, such an overring is guaranteed to exists.
Proof.
We start by noting that for any adjacent domain $D\subseteq D_{1}$ there exists a unique $\mathfrak{p}\in\text{Spec}(D)$ such that $\mathfrak{p}D_{1}=D_{1}$ by Theorem 6.3. Thus, we must have $D_{1}=\cap_{\mathfrak{q}\neq\mathfrak{p}}D_{\mathfrak{q}}=D[\frac{1}{a}]$ where $aD=\mathfrak{q}^{n}$ and $n=|[\mathfrak{q}]|$. Recall that $D[\frac{1}{a}]$ is the localization of $D$ at the complement of the set-theoretic union of all prime ideals of $D$ not containing $a$. In this case, that is all prime ideals omit $\mathfrak{q}$. By continuing this process and recalling the correspondence between prime ideals of a localization $D_{S}$ and prime ideals of $D$ disjoint from $S$ given by $\mathfrak{p}\mapsto\mathfrak{p}D_{S}$, we see that there is a unique, finite set of $r_{1}$ primes $\mathfrak{p}_{1},...,\mathfrak{p}_{r_{1}}\in\text{Spec}(D)$ such that $\mathfrak{p}R=R$. Thus, any sequence of adjacent domains between $D$ and $R$ must have length $r_{1}:=r$ and is unique up to our ordering of the $\mathfrak{p}_{i}$. In particular, if $\mathfrak{p}_{i}D_{i-1}$ is the unique prime such that $\mathfrak{p}_{i}D_{i}=D_{i}$, we must have $D_{i}=D_{i-1}[\frac{1}{a_{i}}]$ where $a_{i}D_{i-1}=\mathfrak{p}_{i}^{n_{i}}$.
Now, to construct such a PID and sequence, take some non-principal $\mathfrak{q}_{1}\in D$ such that $\mathfrak{q}_{1}^{n_{1}}=b_{1}D$, and let $D_{1}=D[\frac{1}{b}]$. As $\mathfrak{q}_{1}]\neq\text{Prin}(D)$, we must have $|Cl(D_{1})|<|Cl(D)|$. By the finiteness of the class group, if we continue this process, we will eventually have $|Cl(D_{t})|=1$ for some $t\in\mathbb{N}$ with $|Cl(D_{i})|\neq 1$ for all $0<i<t$. Thus, $D_{t}$ will be minimal with respect ot being a PID. ∎
This theorem tells us that any ring of integers in particular is finitely many adjacent steps away from being a PID.